# Turb-L1: Achieving Long-term Turbulence Tracing By Tackling Spectral Bias

### v2

@@ Abstract
Accurately predicting the long-term evolution of turbulence is crucial for advancing scientific understanding and optimizing engineering applications. However, existing deep learning methods face significant bottlenecks in long-term autoregressive prediction, which exhibit excessive smoothing and fail to accurately track complex fluid dynamics. Our extensive experimental and spectral analysis of prevailing methods provides an interpretable explanation for this shortcoming, identifying Spectral Bias as the core obstacle. Concretely, spectral bias is the inherent tendency of models to favor low-frequency, smooth features while overlooking critical high-frequency details during training, thus reducing fidelity and causing physical distortions in long-term predictions. Building on this insight, we propose Turb-L1, an innovative turbulence prediction method, which utilizes a Hierarchical Dynamics Synthesis mechanism within a multi-grid architecture to explicitly overcome spectral bias. It accurately captures cross-scale interactions and preserves the fidelity of high-frequency dynamics, enabling reliable long-term tracking of turbulence evolution. Extensive experiments on the 2D turbulence benchmark show that Turb-L1 demonstrates excellent performance: (I) In long-term predictions, it reduces Mean Squared Error (MSE) by $80.3\%$ and increases Structural Similarity (SSIM) by over 9×9$\times$9 × compared to the SOTA baseline, significantly improving prediction fidelity. (II) It effectively overcomes spectral bias, accurately reproducing the full enstrophy spectrum and maintaining physical realism in high-wavenumber regions, thus avoiding the spectral distortions or spurious energy accumulation seen in other methods. (III) It achieves outstanding long-term stability with a relative L2 error as low as $0.444$, substantially outperforming baseline methods and proving its robustness. Our codes are available at https://github.com/easylearningscores/TurbL1_AI4Science.
11footnotetext: These authors contributed equally.22footnotetext: Corresponding authors.
@@ 1 Introduction
Turbulence sreenivasan1999fluid ; wu2024beamvq ; davidson2015turbulence ; duraisamy2019turbulence , the intricate and chaotic dance of fluids shaping phenomena from weather systems to aerodynamic efficiency bi2023accurate ; lam2023learning ; wu2024earthfarsser ; gao2025oneforecast . Accurate prediction of its long-term evolution is paramount not only for fundamental scientific understanding in fields like astrophysics brandenburg2013astrophysical and climate modeling jackson2008parameterization ; wyngaard2010turbulence ; shu2025advanced but also for critical engineering applications such as optimizing energy conversion pope2001turbulent ; borghi1988turbulent and reducing drag kim2003control ; jimenez2018coherent . However, its inherent multiscale nature kolmogorov1995turbulence ; kolmogorov1991local , characterized by strong nonlinearity ottino1989kinematics ; monin1978nature , chaotic dynamics, and the crucial energy cascade across scales, where fine structures dictate macroscopic behavior, renders long-term, high-fidelity prediction an enduring scientific frontier moin1998direct ; duraisamy2019turbulence ; feiopen ; wang2024modeling ; wang2024nuwadynamics .
Traditional numerical simulations for fluid dynamics, such as direct numerical simulation (DNS) moin1998direct ; scardovelli1999direct ; lee2015direct , large eddy simulation (LES) zhiyin2015large ; piomelli1999large , and Reynolds average Navier-Stokes (RANS) models alfonsi2009reynolds ; girimaji2006partially , while foundational, grapple with a persistent trade-off between computational cost and physical accuracy, particularly hindering long-duration, high-resolution turbulence modeling. This challenge spurred interest in deep learning, which, despite successes in tasks like reconstruction thapa2020dynamic ; liu2020deep and short-term forecasting lopez2021model ; li2021fourier ; wu2024earthfarsser , encounters a critical and widespread limitation when applied to the demanding problem of long-term autoregressive turbulence prediction using current mainstream architectures (e.g., Convolutional neural networks (CNNs) he2019mgnet , Transformers wu2024transolver , Neural Operators wu2024neural ; li2021fourier ). Inherent architectural characteristics, such as resolution loss in CNNs, difficulty capturing localized high-frequencies in Transformers, or explicit/implicit filtering in Operators—often lead to a failure to preserve crucial fine-scale dynamics. Consequently, while initial predictions may seem plausible, extended rollouts rapidly diverge from physical reality, manifesting as excessive smoothing, loss of vital vortex structures, severe energy spectrum distortions, and unphysical artifacts, as shown in Figure. 1a, ultimately undermining their reliability for long-range forecasting needs that traditional methods also struggled to meet efficiently.
@@
Figure 1: Long-term turbulence prediction performance comparison. (a) Visual comparison of vorticity fields ($t$=20, $t$=99). Turb-L1 maintains high fidelity to the Ground-truth, while baselines exhibit excessive smoothing (e.g., FNO) or simulation collapse with artifacts (e.g., SimVP) driven by high-frequency error amplification. (b) Enstrophy spectrum density comparison ($t$=10, 30, 99). Baselines fail to capture high-frequency dynamics, increasingly deviating from the Ground Truth at high wavenumbers ($k$) over time, whereas Turb-L1 accurately preserves spectral characteristics.
@@
This study delves into the fundamental reasons underpinning this prevalent failure across diverse deep learning architectures. We identify and argue that the core issue stems from an inherent property of many standard models: spectral bias rahaman2019spectral . This bias manifests as a tendency during training to preferentially learn smooth, large-scale, low-frequency components, while struggling to accurately capture and maintain the less energetic, yet dynamically critical, high-frequency, small-scale structures. For turbulence, whose very essence lies in the intricate cross-scale energy cascade and the dynamics of fine-scale vortices governing dissipation, this inherent learning preference is particularly destructive. It leads models to systematically neglect or misrepresent the behavior of these vital high-frequency components, disrupting the proper energy transfer across scales. As a direct consequence, even small initial errors, particularly in the poorly learned high-frequency range, rapidly amplify during the autoregressive prediction process within the nonlinear system. This amplification directly leads to the observed degradation of fine structures and emergence of artifacts seen visually (Figure.1a), manifests as anomalous high-wavenumber energy behavior and severe spectral distortions (Figure. 1b), and ultimately hinders the ability to accurately trace the turbulence evolution over extended periods.
@@
Addressing the limitations of existing deep learning methods often compromised by spectral bias in long-term turbulence forecasting, this paper presents Turb-L1. Crucially, the design of Turb-L1 is directly motivated by our interpretable identification of spectral bias as the primary impediment to accurate long-term prediction. Our method stems from a more foundational principle: learning the intrinsic spectral dynamics across interacting scales. To this end, we introduce the Hierarchical Dynamics Synthesis mechanism within a multi-grid framework (Figure 2). This process first employs operators highly sensitive to local, high-frequency variations to capture fine-grained turbulent features; subsequently, it utilizes mechanisms adept at global context integration to model the evolution of large-scale structures and their coupling with finer scales. Through the careful orchestration of this two-stage synergistic spectral feature synthesis process, explicit refining of high-frequency components before their integration into the global context, Turb-L1 effectively captures the complex cross-scale energy transfer fundamental to turbulence. As demonstrated visually and spectrally in Figure 1, unlike baseline methods that succumb to excessive smoothing or simulation collapse, our approach preserves high-fidelity details and maintains physical consistency over extended rollouts. This design enables robust and accurate long-range predictions crucial for scientific and engineering applications. We conducted extensive experiments on challenging turbulence benchmarks, where Turb-L1 achieves consistent state-of-the-art (SOTA) performance, demonstrating significant gains in long-term fidelity and stability.
Experimental Observations. Comparative experiments on the McWilliams 2D turbulence benchmark mcwilliams1984emergence validate the effectiveness of Turb-L1. ❶ Delivers state-of-the-art long-term accuracy and fidelity: At $t=99$, Turb-L1 reduces MSE to $0.617$, an $80.3\%$ improvement over the best baseline FNO, and maintains SSIM at $0.682$, over 9×9$\times$9 × higher than baselines like FNO, SimVP, and UNet, successfully preserving fine-scale structures. ❷ Overcomes spectral bias, maintaining physical realism: As shown in Figure 1b, it accurately captures the full enstrophy spectrum, crucially preserving high-frequency dynamics. This avoids issues like the excessive smoothing seen in FNO or the error explosion leading to simulation collapse observed in SimVP, where MSE increased over 14×14$\times$14 × from $t=20$ to $t=99$. ❸ Exhibits superior long-term stability: At $t=99$, the relative L2 error is only $0.444$. This is at least $57\%$ lower than baselines typically exceeding $1.0$, such as FNO at $1.041$ and SimVP at $1.321$. This ensures reliable tracking of turbulence dynamics, highlighting its significant potential for long-term scientific and engineering prediction tasks.
@@ 2.1 Problem Definition
$\frac{\partial\omega(\mathbf{x},t)}{\partial t}+\nabla\cdot(\mathbf{u}(\mathbf{x},t)\omega(\mathbf{x},t))=\nu\nabla^{2}\omega(\mathbf{x},t)+f(\mathbf{x},t)$ (1)
@@
$\hat{\omega}(\mathbf{x},t_{k+1})=\mathcal{F}(\hat{\omega}(\mathbf{x},t_{k});\theta)$ (2)
@@ 2.2 Overview of the Proposed Method: Turb-L1
$\hat{\omega}(\mathbf{x},t_{k+1})=(\mathcal{D}\circ\mathcal{HDS}\circ\mathcal{E})(\omega(\mathbf{x},t_{k});\theta)\equiv\mathcal{F}(\omega(\mathbf{x},t_{k});\theta)$ (3)
@@ 2.3 Projecting Turbulence onto a Latent Manifold Space
Inspired by multi-grid methods, our encoder $\mathcal{E}$ hierarchically projects data onto a latent manifold $\mathcal{M}_{\omega}$ where dynamics are assumed to be smoother. This involves restriction operators $\mathcal{R}_{l}$ and learned transformations $\phi_{l}(\cdot;\theta_{\mathcal{E}}^{(l)})$. The restriction operator $\mathcal{R}_{l}$, as shown in Figure 2’Restriction, showing higher information density to a coarser grid via local aggregation or pooling. This crucial step reduces dimensionality and acts as a localized filter, preferentially passing smoother, large-scale signal components to coarser levels. The learned transformations $\phi_{l}$ then process these restricted features:
@@
$\mathbf{z}^{(l+1)}(\mathbf{x}^{\prime},t_{k})=\phi_{l+1}(\mathcal{R}_{l}(\mathbf{z}^{(l)}(\mathbf{x},t_{k}));\theta_{\mathcal{E}}^{(l+1)}),\quad\text{for }l=0,\dots,L-1$ (5)
where $\mathbf{z}^{(l)}$ is the latent representation at grid level $l$. This multi-grid encoding hierarchically disentangles features for effective cross-scale modeling, enhances computational efficiency, and facilitates dynamics learning on smoother latent manifolds by mitigating the stiffness and high dimensionality of the original PDE solution space, thus providing a rich, multi-scale context for the HDS mechanism.

### v3

@@ Abstract
Accurately predicting the long-term evolution of turbulence is crucial for advancing scientific understanding and optimizing engineering applications. However, existing deep learning methods face significant bottlenecks in long-term autoregressive prediction, which exhibit excessive smoothing and fail to accurately track complex fluid dynamics. Our extensive experimental and spectral analysis of prevailing methods provides an interpretable explanation for this shortcoming, identifying Spectral Bias as the core obstacle. Concretely, spectral bias is the inherent tendency of models to favor low-frequency, smooth features while overlooking critical high-frequency details during training, thus reducing fidelity and causing physical distortions in long-term predictions. Building on this insight, we propose Turb-L1, an innovative turbulence prediction method, which utilizes a Hierarchical Dynamics Synthesis mechanism within a multi-grid architecture to explicitly overcome spectral bias. It accurately captures cross-scale interactions and preserves the fidelity of high-frequency dynamics, enabling reliable long-term tracking of turbulence evolution. Extensive experiments on the 2D turbulence benchmark show that Turb-L1 demonstrates excellent performance: (I) In long-term predictions, it reduces Mean Squared Error (MSE) by $80.3\%$ and increases Structural Similarity (SSIM) by over $9\times$ compared to the SOTA baseline, significantly improving prediction fidelity. (II) It effectively overcomes spectral bias, accurately reproducing the full enstrophy spectrum and maintaining physical realism in high-wavenumber regions, thus avoiding the spectral distortions or spurious energy accumulation seen in other methods. (III) It achieves outstanding long-term stability with a relative $L_{2}$ error as low as $0.444$, substantially outperforming baseline methods and proving its robustness. Our codes are available at https://github.com/Alexander-wu/TurbL1_AI4Science
Index Terms:
Deep Learning, Long-term Prediction, Spectral Bias
@@ 1 Introduction
Turbulence [sreenivasan1999fluid, davidson2015turbulence, duraisamy2019turbulence], the intricate and chaotic dance of fluids shaping phenomena from weather systems to aerodynamic efficiency [bi2023accurate, lam2023learning, wu2024earthfarsser, gao2025oneforecast]. Accurate prediction of its long-term evolution is paramount not only for fundamental scientific understanding in fields like astrophysics [brandenburg2013astrophysical] and climate modeling [jackson2008parameterization, wyngaard2010turbulence] but also for critical engineering applications such as optimizing energy conversion [pope2001turbulent, borghi1988turbulent] and reducing drag [kim2003control, jimenez2018coherent]. However, its inherent multiscale nature [kolmogorov1995turbulence, kolmogorov1991local], characterized by strong nonlinearity [ottino1989kinematics, monin1978nature], chaotic dynamics, and the crucial energy cascade across scales, where fine structures dictate macroscopic behavior, renders long-term, high-fidelity prediction an enduring scientific frontier [moin1998direct, duraisamy2019turbulence].
@@
Figure 1: Long-term turbulence prediction performance comparison. (a) Visual comparison of vorticity fields (tt=20, tt=99). Turb-L1 maintains high fidelity to the Ground-truth, while baselines exhibit excessive smoothing (e.g., FNO) or simulation collapse with artifacts (e.g., SimVP) driven by high-frequency error amplification. (b) Enstrophy spectrum density comparison (tt=10, 30, 99). Baselines fail to capture high-frequency dynamics, increasingly deviating from the Ground Truth at high wavenumbers ($k$) over time, whereas Turb-L1 accurately preserves spectral characteristics.
Traditional numerical simulations for fluid dynamics, such as direct numerical simulation (DNS) [moin1998direct, scardovelli1999direct, lee2015direct], large eddy simulation (LES) [zhiyin2015large, piomelli1999large], and Reynolds average Navier-Stokes (RANS) models [alfonsi2009reynolds, girimaji2006partially], while foundational, grapple with a persistent trade-off between computational cost and physical accuracy, particularly hindering long-duration, high-resolution turbulence modeling. This challenge spurred interest in deep learning, which, despite successes in tasks like reconstruction [thapa2020dynamic, liu2020deep] and short-term forecasting [lopez2021model, li2021fourier, wu2024earthfarsser], encounters a critical and widespread limitation when applied to the demanding problem of long-term autoregressive turbulence prediction using current mainstream architectures (e.g., Convolutional neural networks (CNNs) [he2019mgnet], Graph neural networks (GNNs) [lam2023learning, gao2025oneforecast], Transformers [wu2024transolver], Neural Operators [li2020neural, li2021fourier, raonic2023convolutional, wu2024neural]). Inherent architectural characteristics, such as resolution loss in CNNs, difficulty capturing localized high-frequencies in Transformers, or explicit/implicit filtering in Operators often lead to a failure to preserve crucial fine-scale dynamics. Consequently, while initial predictions may seem plausible, extended rollouts rapidly diverge from physical reality, manifesting as excessive smoothing, loss of vital vortex structures, severe energy spectrum distortions, and unphysical artifacts, as shown in Figure. 1a, ultimately undermining their reliability for long-range forecasting needs that traditional methods also struggled to meet efficiently.
@@
This study delves into the fundamental reasons underpinning this prevalent failure across diverse deep learning architectures. We identify and argue that the core issue stems from an inherent property of many standard models: spectral bias [rahaman2019spectral, fridovich2022spectral, shi2022measuring]. This bias manifests as a tendency during training to preferentially learn smooth, large-scale, low-frequency components [sitzmann2020implicit], while struggling to accurately capture and maintain the less energetic, yet dynamically critical, high-frequency, small-scale structures [tancik2020fourier]. For turbulence, whose very essence lies in the intricate cross-scale energy cascade and the dynamics of fine-scale vortices governing dissipation, this inherent learning preference is particularly destructive [karniadakis2021physics]. It leads models to systematically neglect or misrepresent the behavior of these vital high-frequency components, disrupting the proper energy transfer across scales [bai2021predicting]. As a direct consequence, even small initial errors, particularly in the poorly learned high-frequency range, rapidly amplify during the autoregressive prediction process within the nonlinear system [pathak2022fourcastnet]. This amplification directly leads to the observed degradation of fine structures and emergence of artifacts seen visually (Figure. 1a), manifests as anomalous high-wavenumber energy behavior and severe spectral distortions (Figure. 1b), and ultimately hinders the ability to accurately trace the turbulence evolution over extended periods.
@@
Addressing the limitations of existing deep learning methods often compromised by spectral bias in long-term turbulence forecasting, this paper presents Turb-L1. Crucially, the design of Turb-L1 is directly motivated by our interpretable identification of spectral bias as the primary impediment to accurate long-term prediction. Our method stems from a more foundational principle: learning the intrinsic spectral dynamics across interacting scales. To this end, we introduce the Hierarchical Dynamics Synthesis mechanism within a multi-grid framework. This process first employs operators highly sensitive to local, high-frequency variations to capture fine-grained turbulent features; subsequently, it utilizes mechanisms adept at global context integration to model the evolution of large-scale structures and their coupling with finer scales. Through the careful orchestration of this two-stage synergistic spectral feature synthesis process, explicit refining of high-frequency components before their integration into the global context, Turb-L1 effectively captures the complex cross-scale energy transfer fundamental to turbulence. As demonstrated visually and spectrally in Figure 1, unlike baseline methods that succumb to excessive smoothing or simulation collapse, our approach preserves high-fidelity details and maintains physical consistency over extended rollouts. This design enables robust and accurate long-range predictions crucial for scientific and engineering applications. We conducted extensive experiments on challenging turbulence benchmarks, where Turb-L1 achieves consistent state-of-the-art (SOTA) performance, demonstrating significant gains in long-term fidelity and stability.
Experimental Observations. Comparative experiments on the McWilliams 2D turbulence benchmark [mcwilliams1984emergence] validate the effectiveness of Turb-L1. ❶ Delivers state-of-the-art long-term accuracy and fidelity: At $t=99$, Turb-L1 reduces MSE to $0.617$, an $80.3\%$ improvement over the best baseline FNO, and maintains SSIM at $0.682$, over $9\times$ higher than baselines like FNO, SimVP, and UNet, successfully preserving fine-scale structures. ❷ Overcomes spectral bias, maintaining physical realism: As shown in Figure 1b, it accurately captures the full enstrophy spectrum, crucially preserving high-frequency dynamics. This avoids issues like the excessive smoothing seen in FNO or the error explosion leading to simulation collapse observed in SimVP, where MSE increased over $14\times$ from $t=20$ to $t=99$. ❸ Exhibits superior long-term stability: At $t=99$, the relative L2 error is only $0.444$. This is at least $57\%$ lower than baselines typically exceeding $1.0$, such as FNO at $1.041$ and SimVP at $1.321$. This ensures reliable tracking of turbulence dynamics, highlighting its significant potential for long-term scientific and engineering prediction tasks.
@@
2 Related Work
2.1 Traditional Numerical Simulation of Turbulence
Traditional methods for turbulence simulation, such as Direct Numerical Simulation (DNS) [moin1998direct, scardovelli1999direct], Large Eddy Simulation (LES) [zhiyin2015large, piomelli1999large], and Reynolds-Averaged Navier-Stokes (RANS) [alfonsi2009reynolds, girimaji2006partially], serve as the foundation of fluid dynamics research. While DNS provides the highest fidelity by resolving all spatio-temporal scales, its prohibitive computational cost makes it impractical for long-term simulations at high Reynolds numbers common in engineering applications [lee2015direct]. LES and RANS reduce this cost by modeling sub-grid scales or averaging turbulent fluctuations, respectively. However, these approaches rely on closure models that often lack universality [maulik2017neural] and can introduce significant errors in complex flow regimes [ling2016reynolds, duraisamy2019turbulence, kochkov2021machine]. These methods present a fundamental trade-off between computational efficiency and physical accuracy. In contrast to these methods, which depend on hand-crafted models, our work aims to learn the complete evolution operator directly from high-fidelity data, seeking to achieve near-DNS accuracy in long-term predictions at a fraction of the computational cost.
2.2 Deep Learning for Physical System Modeling
Deep learning has recently shown immense potential for modeling physical systems. Models based on Convolutional Neural Networks (CNNs), such as U-Net [ronneberger2015u] and ResNet [he2016deep], excel at capturing local spatial features but are prone to losing high-frequency information through repeated pooling and convolution operations, leading to over-smoothing in long-term autoregressive predictions [sanchez2020learning, wu2024pure]. The Fourier Neural Operator (FNO) [li2021fourier, kovachki2023neural] learns PDE solution operators efficiently by performing global convolutions in the frequency domain; however, its nature as a low-pass filter systematically neglects small-scale dynamics. Transformer architectures [gao2022earthformer, lam2023learning, wu2024transolver, wu2025advanced] capture global dependencies but are not inherently sensitive to local, high-frequency structural changes. Prevailing deep learning approaches, including CNNs, Transformers, and spatiotemporal models like SimVP [tan2022simvp], universally struggle with either over-smoothing or simulation collapse in long-term turbulence forecasting. This paper systematically argues that the common root of these failures is the inherent spectral bias [rahaman2019spectral, kiessling2022computable, fang2024addressing] of neural networks.
2.3 Architectures for Addressing Spectral Bias
Spectral bias, the tendency of neural networks to preferentially learn low-frequency, smooth functions [rahaman2019spectral], is a key reason for the loss of high-frequency details. Although prior work has leveraged multi-grid concepts (e.g., MGNet [he2019mgnet]) to improve network performance or attempted to implicitly introduce high-frequency information through techniques like positional encodings, these approaches do not fundamentally resolve the issue [mildenhall2021nerf]. Our model, Turb-L1, introduces a novel and explicit solution. It not only employs a multi-grid framework for scale separation but, more critically, incorporates a Hierarchical Dynamics Synthesis (HDS) mechanism. This mechanism forces the model to learn and preserve full-spectrum dynamics by routing features through specialized pathways—using CNNs for high frequencies [dai2021coatnet, wu2021cvt, guo2022cmt] and Transformer-like operators [dosovitskiy2020image, wu2024spatio] for low frequencies. This architecture-driven strategy, designed to directly counteract spectral bias, is the core distinction between Turb-L1 and existing methods and is the key to its unprecedented performance in long-term, high-fidelity turbulence tracing.
@@ 3.1 Problem Definition
$\frac{\partial\omega(\mathbf{x},t)}{\partial t}+\nabla\cdot(\mathbf{u}(\mathbf{x},t)\omega(\mathbf{x},t))=\nu\nabla^{2}\omega(\mathbf{x},t)+f(\mathbf{x},t),$ (1)
@@
$\hat{\omega}(\mathbf{x},t_{k+1})=\mathcal{F}(\hat{\omega}(\mathbf{x},t_{k});\theta),$ (2)
@@ 3.2 Overview of the Proposed Method: Turb-L1
$\hat{\omega}(\mathbf{x},t_{k+1})=(\mathcal{D}\circ\mathcal{HDS}\circ\mathcal{E})(\omega(\mathbf{x},t_{k});\theta)\equiv\mathcal{F}(\omega(\mathbf{x},t_{k});\theta),$ (3)
@@ 3.3 Projecting Turbulence onto a Latent Manifold Space
Inspired by multi-grid methods, our encoder $\mathcal{E}$ hierarchically projects data onto a latent manifold $\mathcal{M}_{\omega}$ where dynamics are assumed to be smoother. This involves restriction operators $\mathcal{R}_{l}$ and learned transformations $\phi_{l}(\cdot;\theta_{\mathcal{E}}^{(l)})$. The restriction operator $\mathcal{R}_{l}$, as shown in Figure 2, projects features to a coarser grid. In our implementation, each restriction operator is a module comprising a $3\times 3$ strided convolution, followed by GroupNorm and a LeakyReLU activation. The strides are strategically chosen from the sequence ‘[1, 2, 1, 2, …]‘ to ensure a gradual downsampling process. This crucial step reduces dimensionality and acts as a localized filter, preferentially passing smoother, large-scale signal components to coarser levels. The learned transformations $\phi_{l}$ then process these restricted features:
@@
$\mathbf{z}^{(l+1)}(\mathbf{x}^{\prime},t_{k})$ $=\phi_{l+1}(\mathcal{R}_{l}(\mathbf{z}^{(l)}(\mathbf{x},t_{k}));\theta_{\mathcal{E}}^{(l+1)}),$ (5)
$\text{for}~~l=0,\dots,L-1$
where $\mathbf{z}^{(l)}$ is the latent representation at grid level $l$. This multi-grid encoding, backed by our specific convolutional design, hierarchically disentangles features for effective cross-scale modeling, enhances computational efficiency, and facilitates dynamics learning on smoother latent manifolds by mitigating the stiffness and high dimensionality of the original PDE solution space, thus providing a rich, multi-scale context for the HDS mechanism.